# A projection method for particle resampling

### v6

@@ 4.1 PETSc test harness
The testing code for this paper is build on the PETSc-PIC framework [13, 35, 36], a recently developed PIC toolkit in the PETSc [6, 11]. The PETSc-PIC framework primarily relies on two modules to drive forward the particle and finite element space. These modules are DMSwarm [30] and DMPLEX [24], respectively. DMSwarm provides a fully parallel solution for particle methods and for particle-mesh methods while DMPLEX provides generic unstructured mesh creation, manipulation and I/O.
@@
The finite element method (FEM) is used to solve the field equations at each timestep. The PETSc FEM framework abstracts the construction of the finite element using the Ciarlet triple [10], consisting of a mesh object (DMPLEX), a finite-dimensional function space (PetscSpace), and a dual space (PetscDualSpace). This is all handled by the PetscFE object and can be customized from the command line. In previous work [13], simple $H^{1}$ finite element spaces have been sufficient in capturing the short timescale linear plasma kinetics. Thus, we will continue the use of these $H^{1}$ spaces in this work.
@@
Particle pushing for the VP system relies on the characteristics of linear hyperbolic Vlasov equation which may be derived by first written a simplified form of the Vlasov equation,
(7) $\frac{\partial f_{\alpha}}{\partial t}+\frac{\partial\vec{x}}{\partial t}\cdot\nabla_{x}f_{\alpha}+\frac{\partial\vec{v}}{\partial t}\cdot\nabla_{v}f_{\alpha}=0$
@@
where $\mathbf{q}=(\mathbf{x},\mathbf{v})$ is the phase space variable and $\mathbf{z}=(\mathbf{v},-q_{e}\mathbf{E}/m)$ is the combined force. The force term $-q_{e}\mathbf{E}/m$ is independent of velocity, and therefore (7) may be written in the conservative form,
@@
Solving the characteristic equations is conducted in PETSc with the TS module using explicit symplectic integrators, a subclass of geometric integrators introduced by Ruth in [37]. In general, for PIC models, explicit integrators are not energy conservative and have a tendency to increase total energy over long time scales through “numerical heating”. In previous work [22, 13, 35], however, explicit symplectic integrators have been shown to achieve exact conservation of mass and momentum, as well as a stable approximate conservation of the system energy. The PETSc TS module contains well-tested implementations for first- to fourth-order symplectic integrators. For full Tokamak models and collisional cases, the TS module also contains a variety of implicit time integrators. These include the recently added discrete gradients method which has been tested on both VP and collisional Landau systems [14]. With these implicit methods, larger time steps can be taken while remaining stable, capturing long time physical phenomena. Furthermore, exact energy conservation has been previously shown using implicit methods [29]. In this work, however, we are interested in capturing the fastest waves in the Landau damping system. Thus, explicit methods are more appropriate and less costly than implicit integrators. We choose a first-order symplectic integrator, called symplectic Euler.
4.2 Particle grids and finite element order
@@
The phase space continuum grids in this work are on regular 90 degree lattice with an option for simple r-refinement in velocity space. Cartesian sub-particle grids are defined in each phase space cell, similar to Lapenta (Figure 1, [27]). For simplicity, the original grid is used for resampling in this work and adaptivity strategies are left for future work. As discussed in §3.1, with periodic boundary conditions in the spatial dimension and natural boundary conditions in velocity dimension we use at least $Q+1$ particles in each dimension in each phase space cell for stability of the pseudoinverse.
@@ 5 Direct remap method with low-order finite elements
Colella et al. developed a method that can be viewed as a version of our algorithm in a finite volume context. In this algorithm, the continuum grid data after particle deposition, is mapped directly to new particles [43, 32, 33]. Cubic splines are used for particle deposition on Cartesian grids. They demonstrate high-order convergence [32], and use phase-space adaptive mesh refinement (AMR) grids [33]. This method should conserve energy because the cubic spines can represent a second order polynomial exactly and the direct remap obviously conserves all moments.
The test harness supports a direct map method in a vertex-centered, finite element context with $Q1$ elements. The grid vertices define the new particles and the domain is doubly periodic (to make $M_{p}$ square), which is immaterial given that density is negligible at the velocity boundary. The salient feature of this construct is that the particle mass matrix is the identity, $\overline{M}_{p}=I$, and the pseudoinverse vanishes. These experiments are intended to observe the qualitative effect of, in some sense, high-order particle mapping.
@@
Figure 2 show the results of this direct map algorithm with the psuedoinverse project on a linear Landau damping, $A=0.01$, problem with a $128\times 256$ particle grid and $V_{max}=6.0$ and results in agree well with amplitude reported in [32] (Figure 3.1). $Q1$ elements are used for all spaces. This direct map method requires one particle per cell and the projection method requires two particles per cell in each dimension to ensure stability of the pseudoinverse. To maintain the same number of particles for all three tests, the projection method uses a $2\times 2$ particle grid per cell and half as many cells in each dimension. First, this data shows again, with $A=0.01$ the efficacy of resampling in that it suppresses the noise observed without resampling. There appears to be some increase in stability with the pseudoinverse technique, however the coarser FE grid may provide smoothing.
@@ 6 Numerical experiments with pseudoinverse resampling
This section investigates this projection resampling method on a linear and a nonlinear Landau damping tests. One issue to be addressed in a PIC method is ensuring a $C^{0}$ electric field, which our $C^{0}$ discretization of the Poisson equation does not provide given that one order of continuity is lost in the gradient of the potential. One can use an $H(div)$ Poisson solver or $C^{1}$ finite elements, but with $C^{0}$ elements we project the field to the vertices and then back to the particles, resulting a $C^{0}$ electric field. This method helps to stabilize our method, is all but required for $Q2$ elements in space. These instabilities could potentially be addressed with a quite start method [38]. Mesh adaptivity in velocity space [3], may act as a type of quite start, and is a subject of future work.
@@ 6.1 Nonlinear Landau damping
The nonlinear Landau damping test case, defined as setting $A=0.5$, has been studied extensively in the literature (Kraus Table 5.1 tabulates several results of previous work [21]). Cheng and Knorr test with a $32\times 128$ cell continuum grid solver and a time step of $\frac{1}{8}$ that is not only quantitatively similar to our results (Figure 4(b)) but qualitatively similar (Figure 4 in [9] and Figure 5 in [23] shows $\gamma_{1}$ measured with the first two peaks, and the fourth peak while the third is low). Figure 3 and 4 show convergence studies on the amplitude of the electric field for time step, resampling rate and particle and continuum grid resolution.
@@
Figure 4: Converge study of nonlinear Landau damping, $A=0.5$
These convergence tests establish parameters for an highly resolved analysis in Figure 4(b) that agrees well with Cheng and Knorr: $\gamma_{1}=-0.281$ as compared to our rate of $-0.278$, and a growth rate of $0.084$ vs our rate of $0.090$. Figure 3(b) shows convergence study in time step where convergence is observed in the rebound region. The modest effect of resampling in this nonlinear case is observed in Figure 4(a), where resampling appears to increase the amplitude in the rebound stage.
@@ 6.2 Linear Landau damping
The linear case of Landau damping, $A=0.0001$, is used here to demonstrate the potential of the projection algorithm. Figure 5 shows the electric field amplitude with a variety of resampling rates with $Q2$ spaces in both real space and velocity space. A damping rate of $\gamma_{1}=-0.15348$ is observed, which agrees with theory reported in [21].
@@
7 Refinement studies and long time stability
[…]
Figure 6 shows a refinement study in the number of particles per cell which does not effect the location of the rebound. We further note that the level of the rebound converges at about $1M$ particles as is shown in Figure 8(b) and Figure 6 box “(b)”. The amplitude of the second rebound ($T=100$ in units of inverse plasma frequency $[\omega_{p}^{-1}]$) is very close for all but the lowest particle count ($16K$) model and this alternating pattern slowly decays leaving coherent and noise-free waves (see detail at $T=500$ Figure 8(c)). However, the $16K$ particle model starts to heat up and continues to grow at longer times ($T=1,000$ data is available in the repository). The $65K$ and $259K$ models remain stable with clean oscillations up to $T=1,000$. The observed change in amplitude of the rebound may be explained by the plasma’s increased ability to absorb wave energy as the particle density is increased. We expect this to converge monotonically but the additional presence of statistical noise in the system, as well as the rebound’s apparent dependence on its state just before rebounding may offer an explanation as to why we do no observe a clean convergence structure.

### v9

@@ 4 Numerical tests and the PETSc test harness
The testing code for this paper is built on the PETSc-PIC framework [Finn2023, pusztay2023landau, PusztayKnepleyAdams2022], a recently developed PIC toolkit in the PETSc [petsc-web-page, PETScManual]. The PETSc-PIC framework primarily relies on two modules to drive forward the particle and finite element space. These modules are DMSwarm [MayKnepley2017] and DMPLEX [LangeMitchellKnepleyGorman2015], respectively. DMSwarm provides a fully parallel solution for particle methods and for particle-mesh methods while DMPLEX provides generic unstructured mesh creation, manipulation and I/O.
@@
The finite element method (FEM) is used to solve the field equations at each timestep. The PETSc FEM framework abstracts the construction of the finite element using the Ciarlet triple [Ciarlet1976], consisting of a mesh object (DMPLEX), a finite-dimensional function space (PetscSpace), and a dual space (PetscDualSpace). This is all handled by the PetscFE object and can be customized from the command line. In previous work [Finn2023], simple $H^{1}$ finite element spaces have been sufficient in capturing the short timescale linear plasma kinetics. Thus, we will continue the use of these $H^{1}$ spaces in this work.
@@
Particle pushing for the VP system relies on the characteristics of linear hyperbolic Vlasov equation which may be derived by first writing a simplified form of the Vlasov equation,
(5) $\frac{\partial f_{\alpha}}{\partial t}+\frac{\partial\mathbf{x}}{\partial t}\cdot\nabla_{x}f_{\alpha}+\frac{\partial\mathbf{v}}{\partial t}\cdot\nabla_{v}f_{\alpha}=0$
@@
where $\mathbf{q}=(\mathbf{x},\mathbf{v})$ is the phase-space variable and $\mathbf{z}=(\mathbf{v},-q_{e}\mathbf{E}/m)$ is the combined force. The force term $q_{e}\mathbf{E}/m$ is independent of velocity, and therefore (5) may be written in the conservative form,
@@
Solving the characteristic equations is conducted in PETSc with the TS module using explicit symplectic integrators, a subclass of geometric integrators introduced by Ruth in [Ruth1983]. In general, for PIC models, explicit integrators are not energy conservative and have a tendency to increase total energy over long time scales through “numerical heating”. In previous work [Kraus2017, Finn2023, pusztay2023landau], however, explicit symplectic integrators have been shown to achieve exact conservation of mass and momentum, as well as a stable approximate conservation of the system energy. The PETSc TS module contains well-tested implementations for first- to fourth-order symplectic integrators. For full Tokamak models and collisional cases, the TS module also contains a variety of implicit time integrators. These include the recently added discrete gradients method which has been tested on both VP and collisional Landau systems [Finn2025]. With these implicit methods, larger time steps can be taken while remaining stable, capturing long time physical phenomena. Furthermore, exact energy conservation has been previously shown using implicit methods [Markidis2011]. In this work, however, we are interested in capturing the fastest waves in the Landau damping system. Thus, explicit methods are more appropriate and less costly than implicit integrators. We choose a first-order symplectic integrator, symplectic Euler.
4.1 Two-stream instability
In this work, we first study the one-dimensional two-stream instability test, often used as a plasma benchmark [BirdsallLangdon]. This test is characterized by an exponential growth in the electric field, meaning the buildup of statistical noise has less impact on the system than other plasma tests. This makes it a useful control for a resampling algorithm. To focus on the electrostatic kinetic effects, we ignore collisional dynamics and reduce the full VML equations to the collisionless, magnetic-free Vlasov-Poisson (VP) system.
The initial system consists of two counter-streaming thermal electron populations with a small perturbation in physical space, described by the distribution,
(10) $f(x,v,t=0)=\frac{1}{2\sqrt{2\pi v_{th}^{2}}}\left(e^{-(v-v_{d})^{2}/2v_{th}^{2}}+e^{-(v+v_{d})^{2}/2v_{th}^{2}}\right)\left(1+A\cos\left(kx\right)\right),$
(11) $(x,v)=\left[0,4\pi/k\right]\times\left[-v_{max},v_{max}\right],$
where $k=0.5$ and $A=0.01$ are the initial perturbation wavenumber and amplitude, $v_{d}=2.0$ is the drift velocity, i.e., the beam spread, $v_{th}=0.5$ is the thermal velocity of the plasma and $v_{max}=10$. In the test, the plasma wave generated by the particles interacts with each counter-streaming population causing particles moving slower than the wave to be accelerated and particles moving faster than the wave to be decelerated. The net deceleration of particles from the two populations leads to a period of linear growth in the electric field strength which becomes unstable due to the initial fluctuations introduced in the distribution function. After the period of electric field growth, resonant particles ($v\approx v_{th}$) become trapped by oscillatory orbits inside potential wells leading to a saturation and flattening in the electric field strength and formation of phase-space vortices.
The evolution of small linear perturbations in the plasma can be described by the dispersion relation,
(12) $1+\sum_{s}\frac{\omega_{p,s}^{2}}{k^{2}}\int\frac{\partial f_{0}/\partial v}{v-\omega/k}\mathrm{d}v=0,$
which arises from the linearization of the Vlasov–Poisson system. If the slope of the distribution function between the beams is positive at the field’s phase velocity, $v_{\phi}=\omega/k$, i.e., $\partial f_{0}/\partial v>0$, the instability appears. From this relation, we can calculate the linear growth rate of the electric field, $\gamma=Im(\omega)$, as our figure of merit in the verification tests. The later stage phenomena, e.g., electric field saturation and vortex formation, become nonlinear in nature and are no longer described by the dispersion relation. Thus, they are not used for quantitative verification in the test. However, they are useful in observing the qualitative effects of the resampling methods on the grid and particles, as will be discussed in later sections.
The linear growth phase is marked in Figure 3 by the red dotted fit line. We consider the period after the fit line ends to be the beginning of the nonlinear phase. It is near this boundary between linear and nonlinear phases that the second and third phase-space plots in Figure 4 are generated.
4.2 Landau damping
As a more thorough test of noise reduction with resampling, we consider the classic one-dimensional plasma test, Landau damping, given by the initial state,
(13) $f(x,v,t=0)=\frac{1}{\sqrt{2\pi}}e^{-v^{2}}\left(1+A\cos\left(kx\right)\right),$
where $k=0.5$, $v_{max}=6$, and we consider three values for the wave amplitude, $A$: $0.5$, $0.01$ and $0.0001$. In this test, we have normalized the thermal velocity, $v_{th}$, to $1$. The Landau damping test is a popular choice for Vlasov benchmarking because it involves a number of purely kinetic effects, such as phase mixing, and it has simple analytical solutions. A detailed description of the analytical solutions to the Landau damping problem can be found in [Finn2023]. The key difference between this test and two-stream instability is the now negative slope of the distribution, $\partial f_{0}/\partial v<0$, at the wave phase velocity, $v_{\phi}$, which leads to the damping effect observed in the field.
The two cases where the wave amplitude, $A$, is $0.01$ and $0.5$ are referred to as the linear and nonlinear landau damping test cases, respectively [KrausThesis, Cheng1976]. A third test case, $A=0.0001$, sits well within the linear regime and is also common. In the linear case, the small field perturbation is damped out at a rate of $\gamma=-0.153$ in favor of a more homogeneous field. However, in collisionless tests the growth of subgrid modes can disrupt the field damping and cause large gradients to develop in the phase-space. These large gradients lead to a sudden regrowth of the field. Previous work [KrausThesis] has show that the inclusion of collisions can remove these subgrid modes, damping the field to machine precision with a continuum code.
In the nonlinear case, $A=0.5$, the field decay reverses much earlier and the general dynamics differ from that of the linear case. The two primary explanations for this earlier field resurgence are the stronger interaction between the potential well created by the electric field which resonates with and accelerates more of the particles, and the increased phase mixing, evident in phase-space diagrams. To fully capture these dynamics, the nonlinearized form of the Vlasov equation must be considered. Significant work was done by Villani and Mouhot in [Villani2009] to analyze the nonlinear Vlasov equation and show that, while the nonlinear dynamics present in this system lead to a weaker initial decay of the electric field, over long enough time scales, the field will damp out, as it does in the linear case. Thus, it is vital to develop the tools necessary to capture the long time evolution of these kinetic plasmas structures. From [KrausThesis] and [Cheng1976], we expect the initial damping rate of the field to be $\gamma_{1}=-0.286$ which quickly turns into a field growth at a rate of $\gamma_{2}=0.086$. We will use these values to verify our tests in later sections.
4.2.1 Continuum grids
@@
The phase-space continuum grids in the implementation are on regular 90 degree lattice with an option for simple r-refinement in velocity space. Cartesian sub-particle grids are defined in each phase-space cell, similar to Lapenta (Figure 1, [Lapenta1995139]). For simplicity, the original grid is used for resampling in this implementation and adaptivity strategies are left for future work. As discussed in §3.1, with periodic boundary conditions in the spatial dimension and natural boundary conditions in velocity dimension we use at least $Q+1$ particles in each dimension in each phase-space cell for stability of the pseudoinverse.
@@ 5 Direct remap method with low-order finite elements
Colella et al. developed a method that can be viewed as a version of our algorithm in a finite volume context. In this algorithm, the continuum grid data after particle deposition, is mapped directly to new particles [Colella2011, Myers2017, Myers16]. Cubic splines are used for particle deposition on Cartesian grids. They demonstrate high-order convergence [Myers2017], and use phase-space adaptive mesh refinement (AMR) grids [Myers16]. This method should conserve energy because the cubic spines can represent a second order polynomials exactly and the direct remap obviously conserves all moments in projecting the grid to particles.
The test harness supports a direct map method in a vertex-centered, finite element context. The grid vertices define the new particles and the domain is doubly periodic (to make $M_{p}$ square), which is immaterial given that density is negligible at the velocity boundary. The salient feature of this construct is that the particle mass matrix is the identity, $\overline{M}_{p}=I$, and the pseudoinverse vanishes.
@@
Figure 2 show the results of this direct map algorithm with the psuedoinverse project on a linear Landau damping, $A=0.01$, problem with a $128\times 256$ particle grid and $V_{max}=6.0$ and results in agree well with amplitude reported in [Myers2017] (Figure 3.1). $Q1$ elements are used for all spaces. This direct map method requires one particle per cell and the projection method requires two particles per cell in each dimension to ensure stability. To maintain the same number of particles for all three tests, the projection method uses a $2\times 2$ particle grid per cell and half as many cells in each dimension. This data shows again, but with $A=0.01$, the efficacy of resampling in that it suppresses the noise observed without resampling.
@@ 6 Numerical experiments with pseudoinverse resampling
This section investigates this projection resampling method on a two-stream instability, a linear Landau damping and a nonlinear Landau damping test. One issue to be addressed in a PIC method is ensuring a $C^{0}$ electric field, which our $C^{0}$ discretization of the Poisson equation does not provide given that one order of continuity is lost in the gradient of the potential. One can use an $H(div)$ Poisson solver or $C^{1}$ finite elements, but with $C^{0}$ elements we project the field to the vertices and then back to the particles, resulting a $C^{0}$ electric field. This method helps to stabilize our method, and is all but required for $Q2$ elements in space. These instabilities could potentially be addressed with a quite start method [SYDORA1999243]. Mesh adaptivity in velocity space [adams2024performance], may act as a type of quite start, and is a subject of future work.
6.1 Two-stream instability
For the two-stream instability test, we consider a $60\times 400$ particle grid with a $2\times 3$ particle grid per cell. This gives a total particle number of $144,000$. Particle and mesh convergence studies were used to calculate this ideal grid which achieves low error and does not require significant computational costs. As discussed in §4.2.1, $Q1$ elements are used for real space and $Q2$ elements for velocity space.
We consider the results of the two-stream instability test both quantitatively and qualitatively. For the quantitative analysis, we compare the linear growth rate of the electric field for a run without resampling and several runs with different resampling periods to the analytic growth rate. Given the initial distribution defined in §4.1, we expect this analytic growth rate of the electric field to be $\gamma=0.20488$. Figure 3 shows the evolution of the max norm of the electric field, with and without resampling. The linear growth phase is marked in Figure 3 with a dashed red line. Table 1 shows the calculated growth rate and the error in the fit for each run. The data shows a general trend of decreasing error when more resampling steps are taken. However, the improvement in accuracy is small. During each resampling step, we also calculate the second moment, kinetic energy, of the plasma before and after the step. In each of the runs, the change in the kinetic energy during the resampling step is $\mathcal{O}(10^{-12})$, which falls within solver tolerances. Thus, the algorithm is conservative when using $Q2$ elements for velocity space.
(a) Linear growth phase and early nonlinear phase.
(b) Entire runtime.
Figure 3: $E_{max}$, Two-stream instability resampling period comparison test.
Resampling Period $\gamma_{fit}$ $\gamma_{error}\;\left[\%\right]$
1.0 0.20013 2.31684
2.0 0.19999 2.38579
4.0 0.19993 2.41687
8.0 0.19993 2.41358
16.0 0.19985 2.45319
32.0 0.19987 2.44742
No Resampling ($\infty$) 0.19987 2.44742
Table 1: Growth rate errors in two-stream instability resampling rate comparison test.
Qualitatively, we consider the effect resampling has on preserving the grid of particles in the system and on the electric field growth in a long-time simulation. Between $T\sim 90$ and $T\sim 110$, the two “stable” phase-space vortices begin to interact and merge into a single vortex, as seen more clearly in Figure 4. Furthermore, after $T=300$, these vortices disappear altogether, replaced by repeated unstable vortex formations and collapses. These periods of change are also noted in Figure 3, where the electric field briefly declines and then proceeds to grows back stronger. The merging of the vortices causes a breakdown of the coherent charge separation that sustains the electrostatic wave, leading to the brief conversion of field energy back to kinetic energy. The inclusion of resampling appears to shift this vortex merging to earlier times in the simulation. It is likely that the resampling algorithm smooths phase-space filaments and disturbs the relative phase between the particles and wave causing the earlier breakdown of the stable vortices [Bernstein1957]. This breakdown of the stable vortices is also accompanied by an increase of the kinetic entropy of the system. However, this effect is not measured here and will be the focus of future work.
Figure 4: Two-stream instability phase-space density evolution for three cases: (top row) no resampling, (middle row) resampling period $32.0$ and (bottom row) resampling period $1.0$.
Figure 5: Two-stream instability phase-space particles.
In Figure 5, the particles are shown in phase-space with constant weights to remove coloring. For this case, a coarsened particle grid with $(60\times 50)\times(2\times 3)$ was used to better show the effect of resampling on grid preservation. As shown in the figure, prior to the resampling step, the phase-space is significantly distorted, with several regions underrepresented by particles or lacking particles altogether. After resampling, we return to a uniform, well-sampled grid. As shown in Figure 4, over long enough time scales, this effect becomes more pronounced and the need for resampling is obvious. Large secondary vortices appear in phase-space at the top and bottom of the velocity space. In runs with no resampling or high resampling periods (top and middle rows of Figure 4), these vortices are largely void of any particles, whereas, in lower resampling period runs (bottom row of Figure 4), these vortices remain filled with particles and are thus well sampled. Thus, we have shown that the inclusion of the pseudoinverse resampling algorithm leads improved accuracy in the linear growth rate and a more consistently well-sampled grid at later simulation times. However, at early times, the effect of the resampling algorithm is minimal in this test. For a more thorough test of resampling methods, we turn to Landau damping.
@@ 6.2 Nonlinear Landau damping
The nonlinear Landau damping test case, defined as setting $A=0.5$, has been studied extensively in the literature (Kraus Table 5.1 tabulates several results of previous work [KrausThesis]). Cheng and Knorr test with a $32\times 128$ cell continuum grid solver and a time step of $\frac{1}{8}$ that is not only quantitatively similar to our results (Figure 7(b)) but qualitatively similar (Figure 4 in [Cheng1976] and Figure 5 in [Kraus_Kormann_Morrison_Sonnendrücker_2017] shows $\gamma_{1}$ measured with the first two peaks, and the fourth peak). Figure 6 and 7 show convergence studies on the amplitude of the electric field for time step, resampling rate and particle and continuum grid resolution.
@@
Figure 7: Convergence study of nonlinear Landau damping, $A=0.5$
These convergence tests establish parameters for an highly resolved analysis in Figure 7(b) that agrees well with Cheng and Knorr: $\gamma_{1}=-0.281$ as compared to our rate of 0.278-0.278, and a growth rate of $0.084$ vs our rate of $0.090$. Figure 6(b) shows convergence study in time step where convergence is observed in the rebound region. The modest effect of resampling in this nonlinear case is observed in Figure 7(a), where resampling appears to increase the amplitude in the rebound stage.
@@ 6.3 Linear Landau damping
The linear case of Landau damping, $A=0.0001$, is used here to demonstrate the potential of the projection algorithm. Figure 8 shows the electric field amplitude with a variety of resampling rates with $Q2$ spaces in both real space and velocity space. A damping rate of $\gamma_{1}=-0.15348$ is observed, which agrees with theory reported in [KrausThesis].
@@
7 Long time stability
Supporting long time, stable PIC simulation is a well known challenge for PIC methods and reducing noise, as we have demonstrated in even short time frames (Figures 2 and 8), is necessary to address this challenge. This section investigates the behavior of our current method in long time simulations with Landau damping (§7.1) and two-stream (§7.2).
7.1 Landau damping
In an idealized, infinitely resolved grid, we expect the linear Landau damping problem to decay indefinitely [Villani2009]. In a discrete system, however, discretization errors create subgrid modes, with wavelengths bounded by the Nyquist mode, $k_{max}=\pi/\Delta x$. Thus, in coarser grids, high frequency velocity structures in the plasma may alias onto lower-wavenumber subgrid modes, which will drive energy back into the field causing the “rebound” effect observed in essentially all collisionless experiments reported in the literature (Fig. 5.3 [KrausThesis]). This effect can be suppressed by reducing the grid size, $\Delta x$, below the Debye length, $\lambda_{D}=\sqrt{1/n^{3}}$ in 1D normalized form, where $n=\int_{\Omega}fdv$ is the total plasma density. Previous work [Birdsall1980, Werner2025], has suggested an ideal range for suppressing subgrid modes is $0.15\leq\lambda_{D}/\Delta_{x}\leq 1.0$. For our system, at the upper limit of this range ($\lambda_{D}=\Delta x$), given a density of $n=4\pi$ and a Debye length of $\lambda_{D}=0.022$, the number of grid cells will be 560 cells in space.
Figure 9 shows a refinement study in the number of particles per cell which does not effect the location of the rebound. We further note that the amplitude of the rebound converges at about $1M$ particles as is shown in Figure 11(b) and Figure 9 box “(b)”. The amplitude of the second rebound (at time $T=100$ in units of inverse plasma frequency $\omega_{p}^{-1}]$) is very close for all but the lowest particle count ($16K$) model and this alternating pattern slowly decays leaving coherent and noise-free waves (see detail at $T=500$ Figure 11(c)). However, the $16K$ particle model starts to heat up and continues to grow at longer times ($T=1,000$ data is available in the repository). The $65K$ and $259K$ models remain stable with clean oscillations up to $T=1,000$. The observed change in amplitude of the rebound may be explained by the plasma’s increased ability to absorb wave energy as the particle density is increased. We expect this to converge monotonically but the additional presence of statistical noise in the system, as well as the rebound’s apparent dependence on its state just before rebounding may offer an explanation as to why we do no observe a clean convergence structure.